1.21 g of ethanol, C2H5OH, was burned in a spirit burner. The heat produced raised the temperature of 400 g of water placed in a beaker above the flame from 17.0 oC to 29.9 oC. (i) Calculate the enthalpy change, in kJ mol-1, for the reaction taking place. (ii) State four reasons why this value is not equal to – 1371 kJ mol-1 which is the data book value for the standard enthalpy of combustion of ethanol.
1
Expert's answer
2020-04-28T12:20:07-0400

I)

C(H2O) = 4.186 J/g °C

Q = 4.186 × 400 × (29.9 - 17) = 21600 J

n(C2H5OH) = 1.21 / 46 = 0.0263 mol

"\Delta\Eta" = 21600 / 0.0263 = 821293 J / mol = 821.3 kJ / mol


II)

  1. Some heat is lost during burning ethanol
  2. Measurement errors
  3. Water lost some heat because of cooling
  4. Some ethanol evaporated and did not burn